Circular No. 8415

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2004et IN NGC 6946 C. J. Stockdale, Marquette University; K. W. Weiler, Naval Research

Laboratory; S. D. Van Dyk, Spitzer Science Center, California Institute of Technology; R. A. Sramek, National Radio Astronomy Observatory; N. Panagia, European Space Agency and Space Telescope Science Institute; and J. M. Marcaide, University of Valencia, report the detection of radio emission from the type-II supernova 2004et (IAUC 8413, 8414) with the Very Large Array in ‘A’ configuration: “Radio-flux densities on Oct. 5.128 UT of 1.33 ± 0.20 mJy at 22.460 GHz (wavelength 1.3 cm) and 0.34 ± 0.06 = mJy at 8.460 GHz (wavelength 3.6 cm) were observed at α = 20^{h}35^{m}25^{s}.36, δ = +60^{o}07′17′′.7 (equinox 2000.0; ± 0′′.1 in each coordinate). This is in close agreement with the reported optical position. On Sept. 30.18, no radio emission was detected at that position, with 3σ upper limits of 1.29 mJy at 22.460 GHz and 0.29 mJy at 8.460 GHz. This is the eighth supernova recorded in NGC 6946 in the modern era, four of which have been radio-detected; NGC 6946 is the most prolific supernovaproducing galaxy known. SN 2004et has clearly been caught during its turn-on phase for the centimeter radio emission, and observations in other wavebands are urged. Radio observations are continuing.”

SUPERNOVA 2004er IN MCG –01-7-24 M. Modjaz, P. Challis, and R. Kirshner, Harvard-Smithsonian Center

for Astrophysics, report that a spectrum (range 350–740 nm) of SN 2004er (cf. IAUC 8412), obtained by M. Calkins on Oct. 5.40 UT with the Mt. Hopkins 1.5-m telescope (+ FAST spectrograph), shows it to be a type-II supernova. The spectrum consists of a blue continuum and P-Cyg lines of hydrogen and helium. Adopting the NASA/IPAC Extragalactic Database recession velocity of 4411 km/s for the host galaxy, the expansion velocity derived from the minimum of the Hβ line is ∼ 12000 km/s.

V1187 SCORPII R. K. Zamanov, Liverpool John Moores University, reports the follow-

ing CCD magnitude estimates (uncertainties ∼ 0.03 mag) obtained with the 2-m Liverpool telescope on La Palma by M. Bode, I. A. Steele, T. J. O’Brien, and himself: Sept. 26.836 UT, B = 15.46, V = 14.38, R = 12.48, I = 12.13; 27.838, 15.61, 14.43, 12.58, 12.22; 28.831, 15.63, 14.43, 12.61, 12.26; 29.829, 15.61, 14.44, 12.63, 12.27.

2004 October 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green